# CORE4D: A 4D Human-Object-Human Interaction Dataset for Collaborative Object REarrangement

### v1

@@ 3.4 Dataset Characteristics
Figure 4: Dataset statistics. (a) shows some object samples from five categories. Bars in (b) indicate when the person is in contact with the object during the entire collaborative object rearrangement interaction process. (c) presents the proportion of collaboration modes in the dataset.
@@ 4.1 Data Split
We construct a training set from a random assortment of real objects, combining their real motions and corresponding synthetic data. We also create two test sets from CORE4D-Real for non-generalization and inner-category generalization studies. Test set S1 includes interactions with training set objects, while S2 features interactions with new objects. CORE4D-Synthetic is not included in the test set, avoiding potential biases from the retargeting algorithm. Details are shown in supplementary material.
@@ 4.2 Human-object Motion Forecasting
Evaluation Metrics. Following existing motion forecasting works [12, 78, 88], we evaluate human joints position error $J_{e}$, object translation error $T_{e}$, object rotation error $R_{e}$, human-object contact accuracy $C_{\text{acc}}$, and penetration rate $P_{r}$. Details are provided in the supplementary material.
@@
Table 2: Quantitative results on human-object motion forecasting.
Test Set Method Human Object Contact
[…]
S1 MDM [73] 170.8 ($\pm$0.9) 136.8 ($\pm$0.1) 10.7 ($\pm$0.1) 84.9 ($\pm$ 0.2) 0.3 ($\pm$ 0.0)
InterDiff [88] 170.8 ($\pm$0.9) 135.1 ($\pm$0.1) 10.2 ($\pm$0.1) 84.9 ($\pm$ 0.2) 0.3 ($\pm$ 0.0)
CAHMP [12] 169.4 ($\pm$0.3) 110.3 ($\pm$0.1) 9.0 ($\pm$0.1) - -
S2 MDM [73] 186.4 ($\pm$0.7) 136.0 ($\pm$0.2) 11.1 ($\pm$0.0) 88.0 ($\pm$ 0.0) 0.3 ($\pm$ 0.0)
InterDiff [88] 186.4 ($\pm$0.7) 133.6 ($\pm$0.2) 10.7 ($\pm$ 0.1) 88.0 ($\pm$ 0.0) 0.3 ($\pm$ 0.0)
CAHMP [12] 170.5 ($\pm$0.3) 112.9 ($\pm$0.1) 9.5 ($\pm$0.1) - -
@@ 4.3 Interaction Synthesis
Evaluation Metrics. Following individual human-object interaction synthesis [40], we evaluate human joint position error $RR.J_{e}$, object vertex position error $RR.V_{e}$, and human-object contact accuracy $C_{\text{acc}}$. The FID score ($FID$) is leveraged to quantitatively assess the naturalness of synthesized results. Details of the metric designs are presented in the supplementary material.
Methods, Results, and Analysis. We utilize two advanced generative models, MDM [73] and OMOMO [40], as baselines. MDM is a one-stage conditional motion diffusion model, while OMOMO is a two-stage approach with hand positions as intermediate results. Quantitative evaluations reveal larger errors in OMOMO when modeling multi-human collaboration compared to individual interaction synthesis by Li et al. [40]. Furthermore, the synthesized results have a higher FID than real motion data, indicating challenges in motion naturalness.
@@ 4.4 Collaboration Retargeting
Table 4: Ablation study. CC denotes contact candidates. D denotes the human pose discriminator. CCU denotes the contact candidate update.
Comparisons Designs Penetration User Preferences
CC D CCU Distance Contact Motion
(mm, $\downarrow$) A/B/Approx.(%, $\uparrow$)
Abl.1 A 6.06 7.8/88.9/3.3 3.3/84.4/12.3
B 2.38
Abl.2 A 5.49 1.2/91.4/7.4 3.2/85.1/11.7
B 2.38
Abl.3 A 2.38 5.0/76.0/19.0 4.0/69.0/27.0
B 2.27
@@
Ablation on Discriminator. In this ablation, as shown in Table 4.Abl.2, we omit the human pose discriminator in the collaboration retargeting. We will randomly choose a candidate from the contact candidates. There are obvious performance drops, demonstrating the critical role of the human pose discriminator in selecting appropriate candidates.
Ablation on Contact Candidate Update. We exclude contact candidate update process in Table 4.Abl.3 experiment. This removal has weakened our method’s ability to search for optimal solutions on objects, resulting in a modest degradation in penetration distance. The user study still exhibited a strong bias, indicating a perceived decline in the plausibility of both contact and motion. This ablation underscores the importance of contact candidate update within our methodology.
@@
Application of CORE4D-Synthetic. Table 5 compares the motion forecasting ability of light-weighted CAHMP [14]. The test set is S2 defined in Section 4.1. We assess the quality of CORE4D-Synthetic by comparing No.A and No.B. No.A even have better performance on object due to enriched spacial relation between human and object in CORE4D-Synthetic. No.C shows the value of the CORE4D-Synthetic by largely improving the performance. Details are in supplementary material.
@@
No Train Set Human Object
Total Real Synthetic $J_{e}$ (mm, $\downarrow$) $T_{e}$ (mm, $\downarrow$) $R_{e}$ (, $\downarrow$)
A 1.0K 0.1K 0.9K 127.7($\pm$ 0.7) 121.7($\pm$ 0.7) 8.04($\pm$ 0.4)
B 1.0K 1.0K 0 127.0($\pm$ 0.8) 120.5 ($\pm$ 0.6) 9.48 ($\pm$ 0.1)
C 5.0K 1.0K 4.0K 116.2($\pm$ 0.3) 112.1($\pm$ 0.2) 6.99($\pm$ 0.1)
@@ 5 Conclusion and Limitations
Limitations. One limitation is that outdoor scenes are not incorporated due to the usage of the mocap system. Another limitation is that our data augmentation strategy currently focuses on adopting collaboration to novel object geometries while excluding human shape diversity. Integrating our retargeting approach with human shape modeling could be an interesting future direction.
@@
The project page of CORE4D is CORE4D Project Page.
@@
  • I. CORE4D-Real Data Capturing Instructions and Costs
  • J. Experiment Configurations and Codes
  • K. URLs of Dataset, Repository, Metadata, DOI, and License
  • L. Dataset Documentation and Intended Uses
  • M. Author Statement
@@ D.2 Participants
As illustrated in Fig. 5, a total of 31 participants, encompassing variations in height, weight, and gender, contributed to the capturing of CORE4D-Real.
@@ Appendix F Evaluation Metrics for Benchmarks
The code of our evaluation metrics is provided in Code Repository.
@@ F.1 Human-object Motion Forecasting
  • We define $J_{e}$ as the average Mean Per Joint Position Error (MPJPE) of the two persons. MPJPE represents the mean per-joint position error of the predicted human joint positions and the ground-truth values.
  • Translation error ($T_{e}$) and rotation error ($R_{e}$) denote the average L2 difference between the predicted object translation vectors and the ground-truth ones, and the average geodesic difference between the estimated object rotation matrices and the ground-truth ones, respectively.
  • Physical metrics: To assess contact fidelity, we detect contacts on the two hands of the two persons for each frame with an empirically designed distance threshold (5 centimeters). We then examine the contact accuracy ($C_{acc}$), which indicates the average percentage of contact detection errors in the predicted motions. Additionally, we examine the object penetration ratio ($P_{r}$) representing the mean percentage of object vertices inside the human meshes.
@@ F.2 Interaction Synthesis
  • $RR.J_{e}$ denotes the average root-relative MPJPE of the two persons. The root-relative MPJPE represents the mean per-joint position error of the predicted human joint positions relative to the human root position and the ground-truth values.
  • $RR.V_{e}$ denotes the average root-relative Mean Per Vertex Position Error (MPVPE) of the two persons. The root-relative MPVPE represents the mean per-vertex position error of the predicted human vertex positions relative to the human root position and the ground-truth values.
@@
  • The Fréchet Inception Distance (FID) quantitatively evaluates the naturalness of synthesized human motions. We first train a feature extractor on CORE4D-Real to encode each human-object-human motion sequence to a 256D feature vector $\bar{f}_{i}$ and acquire the ground-truth human motion feature distribution $\bar{D}$=$\{\bar{f}_{i}\}$. We then replace the motions of the two persons as synthesized ones and obtain another distribution $D$=$\{f_{i}\}$. Eventually, the FID denotes the 2-Wasserstein distance between $\bar{D}$ and $D$. Since CORE4D-Real provides action labels, the feature extractor is supervised-trained by fulfilling the action recognition task. The network structure of the feature extractor is a single-layer Transformer[75]. We provide the code of the feature extractor and pre-trained parameters in Code Repository.
@@ Appendix G Qualitative Results on Benchmarks
Figure 11 and Figure 12 exemplify generated motions for the human-object motion forecasting task and the interaction synthesis task, respectively, where “GT” denotes the ground truth motions, and others are method predictions. Since the baseline methods do not focus on generating hand poses, we replace hand poses in ground truth with flat hands to facilitate fair comparisons. Despite diverse cooperation modes that can be generated, the baseline methods consistently encompass unsatisfactory performances including unnatural collaboration, inter-penetration, and unnatural contact.
@@ H.3 Results
Appendix I CORE4D-Real Data Capturing Instructions and Costs
I.1 Instructions.

### v2

@@ 3.4 Dataset Characteristics
Figure 4: Dataset statistics. (a) shows object samples from six categories. Bars in (b) indicate when the person is in contact with the object during the entire collaborative object rearrangement interaction process. (c) presents the proportion of collaboration modes in the dataset.
@@ 4.1 Data Split
We construct a training set from a random assortment of real objects, combining their real motions and corresponding synthetic data. We also create two test sets from CORE4D-Real for non-generalization and inner-category generalization studies. Test set S1 includes interactions with training set objects, while S2 features interactions with new objects. CORE4D-Synthetic is not included in the test set, avoiding potential biases from the retargeting algorithm. Details are shown in the appendix.
@@ 4.2 Human-object Motion Forecasting
Evaluation Metrics. Following existing motion forecasting works [14, 102, 119], we evaluate human joints position error $J_{e}$, object translation error $T_{e}$, object rotation error $R_{e}$, human-object contact accuracy $C_{\text{acc}}$, and penetration rate $P_{r}$. Details are provided in the appendix.
@@
Test Method Human Object Contact
Set $J_{e}$ ($\downarrow$) $T_{e}$ ($\downarrow$) $R_{e}$ ($\downarrow$) $C_{\text{acc}}$ ($\uparrow$) $P_{r}$ ($\downarrow$)
S1 MDM [97] 170.8 136.8 10.7 84.9 0.3
InterDiff [119] 170.8 135.1 10.2 84.9 0.3
CAHMP [14] 169.4 110.3 9.0 - -
S2 MDM [97] 186.4 136.0 11.1 88.0 0.3
InterDiff [119] 186.4 133.6 10.7 88.0 0.3
CAHMP [14] 170.5 112.9 9.5 - -
Table 2: Quantitative results on motion forecasting.
@@ 4.3 Interaction Synthesis
Evaluation Metrics. Following individual human-object interaction synthesis [53], we evaluate human joint position error $R.J_{e}$, object vertex position error $R.V_{e}$, and human-object contact accuracy $C_{\text{acc}}$. The FID score (FID) is leveraged to quantitatively assess the naturalness of synthesized results. Details of the metric designs are presented in the appendix.
Methods, Results, and Analysis. We utilize three advanced diffusion models [97, 53, 52] as baselines. MDM [97] and CHOIS [52] are one-stage conditional motion diffusion models, while OMOMO is a two-stage approach with hand positions as intermediate results. Quantitative evaluations reveal larger errors in OMOMO when modeling multi-human collaboration compared to individual interaction synthesis by Li et al. [53]. Furthermore, the synthesized results have a higher FID than real motion data, indicating challenges in motion naturalness.
Test Set Method $R.J_{e}$ ($\downarrow$) $R.V_{e}$ ($\downarrow$) $C_{\text{acc}}$ ($\uparrow$) FID ($\downarrow$)
S1 MDM [97] 138.3 194.8 76.5 7.5
OMOMO [53] 138.0 196.9 78.0 7.8
CHOIS [52] 138.4 194.3 76.2 7.7
S2 MDM [97] 146.1 208.3 76.6 7.9
OMOMO [53] 145.3 209.9 77.8 7.4
CHOIS [52] 145.8 206.7 76.2 7.7
@@ 4.4 Collaboration Retargeting
Comparisons Designs Phys. Eval. User Preferences
C D U $P$ $C_{\text{acc}}$ Contact Motion
($\downarrow$) ($\uparrow$) A / B / Approx. ($\uparrow$)
Abl.1 A 0.61 83.2 7.8/88.9/3.3 3.3/84.4/12.3
B 0.24 83.3
Abl.2 A 0.55 82.9 1.2/91.4/7.4 3.2/85.1/11.7
B 0.24 83.3
Abl.3 A * 0.68 94.7 5.6/84.3/10.1 2.2/86.6/11.2
B 0.24 83.3
Abl.4 A 0.24 83.3 5.0/76.0/19.0 4.0/69.0/27.0
B 0.23 85.5
Table 4: Ablation study. C, D, and U denote contact candidates, the human pose discriminator, and the contact candidate update, respectively. $P$ is penetration distance. $C_{acc}$ is contact accuracy.
@@ C D U $P$ $C_{\text{acc}}$ Contact Motion
Ablations on Discriminator. In Table 4.Abl.2, we omit the human pose discriminator in the collaboration retargeting. Method A randomly chooses a candidate from the contact candidates. There are obvious performance drops, demonstrating the critical role of the human pose discriminator in selecting appropriate candidates. Table 4.Abl.3 further compares the proposed discriminator against selecting the motion with the most accurate contact (method A), and user studies reveal significant superiority of the discriminator.
Ablation on Contact Candidate Update. We exclude contact candidate update process in Table 4.Abl.4 experiment. This removal has weakened our method’s ability to search for optimal solutions on objects, resulting in a modest degradation in penetration distance. The user study still exhibited a strong bias, indicating a perceived decline in the plausibility of both contact and motion. This ablation underscores the importance of contact candidate updates within our methodology.
@@
5 Dataset Applications
5.1 CORE4D-Synthetic Enhances Human-object Motion Forecasting Quality
Table 5 compares the motion forecasting ability of light-weighted CAHMP [16]. The test set is S2 defined in Section 4.1. We assess the quality of CORE4D-Synthetic by comparing No.A and No.B. No.A even have better performance on object due to enriched spacial relation between human and object in CORE4D-Synthetic. No.C shows the value of the CORE4D-Synthetic by largely improving the performance. Details are in the appendix.
No Train Set Human Object
Total Real Synthetic $J_{e}$ ($\downarrow$) $T_{e}$ ($\downarrow$) $R_{e}$ ($\downarrow$)
A 1.0K 0.1K 0.9K 127.7 121.7 8.04
B 1.0K 1.0K 0 127.0 120.5 9.48
C 5.0K 1.0K 4.0K 116.2 112.1 6.99
@@ 5.1 CORE4D-Synthetic Enhances Human-object Motion Forecasting Quality
5.2 CORE4D Supports Humanoid Skill Learning
Benefitting from rapid developments of humanoid robots [25, 47, 100], tremendous progress has been made in studying versatile humanoid skills for locomotion [35, 94, 84, 66] and humanoid-object interaction [68, 21, 133]. Aiming at enabling humanoid learning skills from human data, we select human interaction motions with three large boxes from CORE4D, and retarget them onto the Unitree H1 humanoid robot [100] with object scale augmentation. With 890 human-like humanoid-box interaction sequences, we design a box-lifting task in Isaac Gym [64], and benchmark two state-of-the-art humanoid imitation learning (IL) methodologies [143, 26] comparing to demonstration-free reinforcement learning (RL) paradigm [87].
Table 6 compares the success rates of these methods. Leveraging interaction data from CORE4D, the two IL methods [143, 26] consistently make it possible for humanoids to lift unseen boxes with visual sensor signals successfully, demonstrating that CORE4D can promote humanoid interaction skill learning. Figure 5 exemplifies a successful case of HumanPlus [26], showing that humanoids can learn from CORE4D and achieve the task in a human-like manner. As the development of multi-humanoid imitation learning methods in the future, we anticipate that CORE4D can further promote collaboration skill learning. Details on task formulation, method designs, and evaluations are in the appendix.
Data N/A CORE4D
Method PPO [87] HumanPlus [26] HST [26] + ACT [143]
$SR$ ($\uparrow$) 0.0 21.0 26.5
Table 6: Success rates of RL and IL in humanoid box lifting.
Figure 5: Visualization of the humanoid box-lifting skill trained by CORE4D via imitation learning.
@@ 6 Conclusion and Limitations
Limitations. Firstly, outdoor scenes are not incorporated in CORE4D-Real due to the usage of the mocap system. Secondly, visual signals are excluded in CORE4D-Synthetic. Transferring real-world videos onto synthesized collaboration motions could be an interesting future direction.
@@
  • I. Details on Humanoid Skill Learning using CORE4D
  • J. CORE4D-Real Data Capturing Instructions
@@ D.2 Participants
31 participants, encompassing variations in height, weight, and gender, contributed to the capturing of CORE4D-Real.
@@ F.1 Human-object Motion Forecasting
  • We define $J_{e}$ as the average Mean Per Joint Position Error (MPJPE) of the two persons. MPJPE represents the mean per-joint position error of the predicted human joint positions and the ground-truth values. The unit of $J_{e}$ is one millimeter.
  • Translation error ($T_{e}$) and rotation error ($R_{e}$) denote the average L2 difference between the predicted object translation vectors and the ground-truth ones, and the average geodesic difference between the estimated object rotation matrices and the ground-truth ones, respectively. The unit of $T_{e}$ is one millimeter. The unit of $R_{e}$ is one degree.
  • Physical metrics: To assess contact fidelity, we detect contacts on the two hands of the two persons for each frame with an empirically designed distance threshold (5 centimeters). We then examine the contact accuracy ($C_{acc}$), which indicates the average percentage of contact detection errors in the predicted motions. Additionally, we examine the object penetration ratio ($P_{r}$) representing the mean percentage of object vertices inside the human meshes. The units of the two metrics are percentages.
@@ F.2 Interaction Synthesis
  • $RR.J_{e}$ denotes the average root-relative MPJPE of the two persons. The root-relative MPJPE represents the mean per-joint position error of the predicted human joint positions relative to the human root position and the ground-truth values. The unit of $RR.J_{e}$ is one millimeter.
  • $RR.V_{e}$ denotes the average root-relative Mean Per Vertex Position Error (MPVPE) of the two persons. The root-relative MPVPE represents the mean per-vertex position error of the predicted human vertex positions relative to the human root position and the ground-truth values. The unit of $RR.V_{e}$ is one millimeter.
@@
  • The Fréchet Inception Distance (FID) quantitatively evaluates the naturalness of synthesized human motions. We first train a feature extractor on CORE4D-Real to encode each human-object-human motion sequence to a 256D feature vector $\bar{f}_{i}$ and acquire the ground-truth human motion feature distribution $\bar{D}$=$\{\bar{f}_{i}\}$. We then replace the motions of the two persons as synthesized ones and obtain another distribution $D$=$\{f_{i}\}$. Eventually, the FID denotes the 2-Wasserstein distance between $\bar{D}$ and $D$. Since CORE4D-Real provides action labels, the feature extractor is supervised-trained by fulfilling the action recognition task. The network structure of the feature extractor is a single-layer Transformer [101].
@@ Appendix G Qualitative Results on Benchmarks
Figure 11 and Figure 12 exemplify generated motions for the human-object motion forecasting task and the interaction synthesis task, respectively. Since the baseline methods do not focus on generating hand poses, we replace hand poses in ground truth with flat hands to facilitate fair comparisons. Despite diverse cooperation modes that can be generated, the baseline methods consistently encompass unsatisfactory performances including unnatural collaboration, inter-penetration, and unnatural contact.
@@ H.3 Results
Appendix I Details on Humanoid Skill Learning using CORE4D
As introduced in Section 5.2 in the main paper, we use CORE4D’s human-box interaction data to facilitate humanoid skill learning for box lifting. This section presents details on this application, including the simulation environment configuration (Section I.1), task formulation and evaluation (Section I.2), adapting human interaction data to humanoid (Section I.3I.4), benchmark method designs (Section I.5), and experiments (Section I.6).
I.1 Environment Setup
We use the popular Unitree H1 humanoid robot [100] in Isaac Gym [64] simulation environment. The H1 humanoid has 19 revolute joints with fixed limits on motion ranges. The interaction scene contains the humanoid, a box weighing 0.5kg initially posed on the floor, and a third-person-view camera providing visual signals for the skill policy.
Data clipping: The original interactions in CORE4D are collaborations of two persons. In this application, we regard each of them as two individual human-object interaction motions and obtain the box-lifting clips automatically by measuring hand-object distances and the object’s height. We discard the motion if the individual is not touching the object during its lifting process.
Train-test split: As a result, we acquire 890 individual human-box lifting data clips covering 22 boxes augmented from three real-world ones. We select 16 boxes (504 clips) as the training set and the remains as the test set. The training set is used only to train the skill policy (Section I.5) and its preceding motion tracker (Section I.4), while the test set is used only to evaluate the skill policy to assess its generalizability to unseen box shapes.
Observations: The real-time input of the humanoid skill policy is =$\mathcal{I}=\{s_{\text{proprio}},r_{\text{root}},t_{\text{root}},\mathcal{C},\mathcal{D}\}$, where $s_{\text{proprio}}\in\mathbb{R}^{19\times 2}$ is humanoid’s proprioception, $r_{\text{root}}\in\mathbb{R}^{3}$ and $t_{\text{root}}\in\mathbb{R}^{3}$ are humanoid’s root orientation and position in the world coordinate system respectively, and $\mathcal{C}\in\mathbb{R}^{360\times 480\times 3}$ and $\mathcal{D}\in\mathbb{R}^{360\times 480}$ are the color and depth image captured by the camera. Specifically, $s_{\text{proprio}}$ is composed of the angles and velocities of each joint.
Action: Given the real-time input $\mathcal{I}$, the skill policy needs to generate an action $\mathcal{A}$ and use it to actuate the humanoid in the simulation environment. The action $\mathcal{A}\in\mathbb{R}^{19}$ is defined as 19 DoF joint angles, which are transferred to joint torques via a PD controller with pre-defined proportional-derivative gains [26]. The simulation environment uses the computed joint torques to actuate the humanoid.
I.2 Task Formulation and Evaluation Metric
Given an unseen box in the scene and a starting position, the humanoid is required to adjust its pose, touch the box, and finally lift it larger than 20 centimeters. The evaluation metric (SR) is the task success rate defined as whether the box reaches 20 centimeters higher than its initial position.
I.3 Retargeting Interactions from CORE4D onto H1 Humanoid Robot
Following existing humanoid skill learning advances [35, 36, 94], we use an optimization strategy to solve the retargeting problem. The optimization is finding the optimal sequence of $\{s_{\text{proprio}},r_{\text{root}},t_{\text{root}}\}$ that can minimize differences between motions of the human $A$ and the humanoid $B$ on positions of $K$ paired joints in the world coordinate system $L_{p}=\sum\limits_{k=1}^{K}\|P_{a_{k}}-P_{b_{k}}\|^{2}$, where $P\in\mathbb{R}^{T\times 3}$ denotes the position sequence of a joint, and $T$ is the frame number. $<a_{k},b_{k}>$ is a pre-defined pair of joints with similar semantics, where $a_{k}$ is from the human skeleton, and $b_{k}$ is from the humanoid skeleton.
We design a temporal loss $L_{t}=\sum\limits_{t=2}^{T}\sum\limits_{i=1}^{19}(\|s_{\text{proprio},t,i}-s_{\text{proprio},t-1,i}\|^{2}+\|r_{\text{root},t,i}-r_{\text{root},t-1,i}\|^{2}+\|t_{\text{root},t,i}-t_{\text{root},t-1,i}\|^{2})$ to further improve motion smoothness. The overall optimization target is $L_{p}+L_{t}$.
As a result, this retargeting method converts human-box interaction motions to humanoid-box interaction animations. The animations are not physically realistic, and we use the upcoming tracking method to obtain physically realistic ones following HumanPlus [26].
I.4 Tracking Humanoid-box Interaction Animations
Given a humanoid-box interaction animation, the motion tracking methods [82, 62, 96] control the humanoid in the simulation environment and generate the motion that mostly closely resembles the animation. The generated motions are both physically realistic and able to fulfill the task via mimicking animations, which are utilized as skill demonstrations in the following skill policy learning.
We select HST [26] as our tracker. The official implementation of the HST [26] tracker fails to lift the box, and we design an improved HST that can successfully track the animations and control the humanoid to lift the box physically. We describe the key designs of our tracker below.
Input: In each frame, the tracker inputs the states of the real-time humanoid $S_{h}$, the humanoid animation target $\hat{S_{h}}$, the real-time box $S_{b}$, and the box target $\hat{S_{b}}$. $S_{h}=\{r_{\text{root}},\dot{r}_{\text{root}},t_{\text{root}},\dot{t}_{\text{root}},J,\dot{J},R,\dot{R},T,\dot{T}\}$, where $r_{\text{root}}\in\mathbb{R}^{3}$, $\dot{r}_{\text{root}}\in\mathbb{R}^{3}$, $t_{\text{root}}\in\mathbb{R}^{3}$, $\dot{t}_{\text{root}}\in\mathbb{R}^{3}$, $J\in\mathbb{R}^{19}$, $\dot{J}\in\mathbb{R}^{19}$, $R\in\mathbb{R}^{(19+1)\times 3}$, $\dot{R}\in\mathbb{R}^{(19+1)\times 3}$, $T\in\mathbb{R}^{(19+1)\times 3}$, $\dot{T}\in\mathbb{R}^{(19+1)\times 3}$ denotes root’s global orientation, root’s global angular velocity, root’s global position, root’s global linear velocity, joint angle, joint velocity, global positions of the root and joints, global linear velocities of the root and joints, global orientations of the root and joints, and global angular velocities of the root and joints. $\hat{S_{h}}$ is formulated the same as $S_{h}$. $S_{b}=\{r_{b},\hat{r}_{b},t_{b},\hat{t}_{b}\}$, where $r_{b}$, $\hat{r}_{b}$, $t_{b}$, $\hat{t}_{b}$ denotes box global orientation, box global angular velocity, box global position, and box global linear velocity, respectively. $\hat{S_{b}}$ shares the same definition with $S_{b}$.
Output: In each frame, the tracker outputs a 19 DoF action vector representing target joint angles. The joint torques are computed via a low-level PD controller and are fed into the simulation environment to actuate the humanoid.
Reward function: The tracker is trained via a reinforcement learning (RL) method PPO [87]. The reward design is the most crucial part of the method’s performance. Using original rewards from HumanPlus [26] cannot lift the box successfully due to inaccurate tracked hand positions. To handle this issue, we draw inspiration from PhysHOI [105] and use a multiplication of humanoid reward $\mathcal{R}_{h}$, box reward $\mathcal{R}_{b}$, and humanoid-box interaction reward $\mathcal{R}_{i}$ as the overall reward $\mathcal{R}_{\text{overall}}$: $\mathcal{R}_{\text{overall}}=\mathcal{R}_{h}\times\mathcal{R}_{b}\times\mathcal{R}_{i}$.
  • $\mathcal{R}_{h}=0.5\exp(-5\|T-\hat{T}\|_{1})+5\exp(-10\|t_{\text{root}}-\hat{t}_{\text{root}}\|_{1})+10\exp(-10\|J_{l}-\hat{J}_{l}\|_{1})+10\exp(-10\|J_{r}-\hat{J}_{r}\|_{1})$, where $l$/$r$ denotes the left/right hand.
  • $\mathcal{R}_{b}=\exp(-10\|t_{b}-\hat{t}_{b}\|_{1})$.
  • $\mathcal{R}_{i}=\exp(-10\|P_{l,h\rightarrow b}-\hat{P}_{l,h\rightarrow b}\|_{1}-10\|P_{r,h\rightarrow b}-\hat{P}_{r,h\rightarrow b}\|_{1})$, where $P_{h\rightarrow b}$ and $\hat{P}_{h\rightarrow b}$ denotes the humanoid joint position in the box’s coordinate system, and $l$/$r$ denotes the left/right hand.
Training strategies: We adopt an early-termination strategy from DeepMimic [82] that terminates the rollout when the humanoid is 0.5 meters away from its target or its root height is below 0.5 meters.
I.5 Reinforcement Learning and Imitation Learning Method Designs
Combining the retargeting method (Section I.3) with the improved HST tracker (Section I.4), we transfer CORE4D’s data to physically realistic humanoid box-lifting demonstrations. The final step is to train a skill policy that mimics the demonstrations and can lift unseen boxes in test time. We select two vision-based imitation learning (IL) methods, HIT [26] and ACT [143], and use their official implementations.
To examine the value of the demonstrations, we compare the two IL methods with a commonly used state-based RL algorithm PPO [87]. The PPO is implemented with the code from HumanPlus [26], with a change on the reward design $\mathcal{R}$: $\mathcal{R}=\mathcal{R}_{b}+\mathcal{R}_{\textbf{success}}+\mathcal{R}_{i}+\mathcal{R_{\textbf{alive}}}$, where:
  • $\mathcal{R}_{b}=\exp(\|t_{b}-\hat{t}_{b}\|_{2}^{2})$, where $\hat{t}_{b}$ is the pre-defined target box center position for the task.
  • $\mathcal{R}_{\textbf{success}}=[\mathcal{R}_{d}<0.01]$ encouraging achieving the task.
  • $\mathcal{R}_{i}=-0.1(\|P_{l,h\rightarrow b}\|_{2}^{2}+\|P_{r,h\rightarrow b}\|_{2}^{2})$, where $P_{h\rightarrow b}$ and $\hat{P}_{h\rightarrow b}$ denotes the humanoid joint position in the box’s coordinate system, and $l$/$r$ denotes the left/right hand. This reward encourages humanoid hands to explore near the object.
  • $\mathcal{R_{\textbf{alive}}}=0.1$ encouraging the humanoid being alive.
I.6 Experiments
[…]
Appendix J CORE4D-Real Data Capturing Instructions